OpenStudy (anonymous):

Solve the quadratic equation x^2 − 56 = 10^x by factoring. *Free Medals*

OpenStudy (amoodarya):

\[x^2 − 56 = 10^x \] it is what you type ?

OpenStudy (anonymous):

Yes

OpenStudy (amoodarya):

or \[x^2 − 56 = 10 x \]

OpenStudy (amoodarya):

\[x^2 − 56 = 10^x \] it is not a quadratic equation !

OpenStudy (anonymous):

Solve the quadratic equation x^2 − 56 = 10x by factoring.

OpenStudy (amoodarya):

\[x^2 − 56 = 10x\\x^2-10x − 56 = 0 \\(x-5)^2-5^2-56=0\\(x-5)^2=25+56=81\\(x-2)^2=9^2\\(x-2)=\pm3 \]

OpenStudy (anonymous):

So what's the answer

OpenStudy (amoodarya):

you can find x from that can you go on ?

OpenStudy (anonymous):

I already solved it